# A Cascaded Dilated Convolution Approach for Mpox Lesion Classification

### v3

@@ I Introduction
    I demonstrate that by integrating the novel CAGA module with the EfficientViT backbone [20], the proposed model achieves state-of-the-art (SOTA) accuracy across all evaluated datasets while also maintaining impressive computational efficiency. To assess the effectiveness of this model, I conducted extensive evaluations on three diverse datasets: Mpox Close Skin Image Dataset (MCSI)[17], Monkeypox Skin Images Dataset (MSID)[21], and Monkeypox Skin Lesion Dataset (MSLD)[22][23]. As shown in Fig. 1, the proposed EfficientViT-CAGA model not only achieves the highest accuracy across all datasets but also does so while requiring fewer parameters and FLOPs compared to most other state-of-the-art models.
@@ II-A Vision Transformers
ViT was one of the first pure Transformer-based models to achieve state-of-the-art results on ImageNet. Unlike CNNs, which rely on local receptive fields and other inductive biases, ViT processes images as sequences of patches, enabling a more global understanding of the visual content. However, their lack of inherent inductive biases, such as locality, makes them computationally intensive and heavily dependent on large-scale data[31][32][33][34]. Hybrid architectures[35][36][37][38][39] seek to balance these trade-offs, integrating convolutional operations to model local features while leveraging attention mechanisms for global context. This study employs a hybrid architecture, combining the convolutional backbone of EfficientViT with the Cascaded Atrous Group Attention mechanism for self-attention, resulting in significantly improved computational efficiency and performance.
@@ II-B Deep Learning for Mpox Classification
Figure 2: Integration of CAGA within the Network Architecture (left) and detailed structure of Cascaded Group Attention with Cascaded Atrous Attention Module (right). Left: The MBConv layer represents the Mobile Inverted Bottleneck Convolution[42], which is part of the EfficientViT architecture. A Depthwise Separable Convolution[28] is used to reduce computational complexity while preserving spatial and channel-wise feature interactions, converting the input to an $n\times h$ dimension. Batch Normalization (BN)[43] is applied to the output of the residual connection.
@@ IV Experimental Results and Analysis
This section introduces the dataset and implementation details, followed by a comparison of the proposed approach with state-of-the-art architectures. The proposed methodology is compared to six models, which are shown in Table III. All models mentioned in Table III are pre-trained on ImageNet, except for the classifier head which was randomly initialized following common practice. In contrast, for EfficientViT-CAGA, only the EfficientViT convolutional backbone is pre-trained, while the DSConv and CAGA module are initialized using the Xavier uniform method.
@@ IV-C1 MCSI
Table II reports the comparison of EfficientViT-CAGA with other state-of-the-art models. EfficientViT-CAGA achieves remarkable accuracy while being computationally efficient, with only 4.86G FLOPs, which is 3.5$\times$ less than ViT-Base. Compared to the backbone EfficientViT, proposed EfficientViT-CAGA achieves a significant 37.5% reduction in total parameters while improving performance across all evaluation metrics.

### v4

@@ I Introduction
    I demonstrate that by integrating the novel CAGA module with the EfficientViT-L1 backbone [20], the proposed model achieves state-of-the-art (SOTA) accuracy across all evaluated datasets while also maintaining impressive computational efficiency. To assess the effectiveness of this model, I conducted extensive evaluations on three diverse datasets: Mpox Close Skin Image Dataset (MCSI)[17], Monkeypox Skin Images Dataset (MSID)[21], and Monkeypox Skin Lesion Dataset (MSLD)[22][23]. As shown in Fig. 1, the proposed EfficientViT-CAGA model not only achieves the highest accuracy across all datasets but also does so while requiring fewer parameters and FLOPs compared to most other state-of-the-art models.
@@ II-A Vision Transformers
ViT was one of the first pure Transformer-based models to achieve state-of-the-art results on ImageNet. Unlike CNNs, which rely on local receptive fields and other inductive biases, ViT processes images as sequences of patches, enabling a more global understanding of the visual content. However, their lack of inherent inductive biases, such as locality, makes them computationally intensive and heavily dependent on large-scale data[31][32][33][34]. Hybrid architectures[35][36][37][38][39] seek to balance these trade-offs, integrating convolutional operations to model local features while leveraging attention mechanisms for global context. This study employs a hybrid architecture, combining the convolutional backbone of EfficientViT-L1 with the Cascaded Atrous Group Attention mechanism for self-attention, resulting in significantly improved computational efficiency and performance.
@@ II-B Deep Learning for Mpox Classification
Figure 2: Integration of CAGA within the Network Architecture (left) and detailed structure of Cascaded Group Attention with Cascaded Atrous Attention Module (right). Left: The MBConv layer represents the Mobile Inverted Bottleneck Convolution[42], which is part of the EfficientViT-L1 architecture. A Depthwise Separable Convolution[28] is used to reduce computational complexity while preserving spatial and channel-wise feature interactions, converting the input to an $n\times h$ dimension. Batch Normalization (BN)[43] is applied to the output of the residual connection.
@@ IV Experimental Results and Analysis
This section introduces the dataset and implementation details, followed by a comparison of the proposed approach with state-of-the-art architectures. The proposed methodology is compared to six models, which are shown in Table III. All models mentioned in Table III are pre-trained on ImageNet, except for the classifier head which was randomly initialized following common practice. In contrast, for EfficientViT-CAGA, only the EfficientViT-L1 convolutional backbone is pre-trained, while the DSConv and CAGA module are initialized using the Xavier uniform method.
@@ IV-C1 MCSI
Table II reports the comparison of EfficientViT-CAGA with other state-of-the-art models. EfficientViT-CAGA achieves remarkable accuracy while being computationally efficient, with only 4.86G FLOPs, which is 3.5$\times$ less than ViT-Base. Compared to the backbone EfficientViT-L1, proposed EfficientViT-CAGA achieves a significant 37.5% reduction in total parameters while improving performance across all evaluation metrics.